Determine the H*. OH , and pOH of a solution with a pH of 9.57 at 25 °C. [OH] pOH
Answer #1

we have

pH = 9.57

We know the formulas

pH = -log[H+]

pOH = -log[OH-]

pH + pOH = 14

Now

pH = 9.57 = -log[H+]

log[H+] = -pH = -9.57

hence

[H+] = 10-pH = 10-9.57 =

So

[H+] = 2.69 \times 10-10 M

now

pH + pOH = 14

hence

pOH = 14 - pH = 14 - 9.57 = 4.43

pOH = 4.43

Now

pOH = -log[OH-]

log[OH-] = -pOH

hence

[OH-] = 10-pOH = 10-4.43 =

[OH-] = 3.71 \times 10-5 M

So answers are

[H+] = 2.69 \times 10-10 M

[OH-] = 3.71 \times 10-5 M

pOH = 4.43